OpenStudy (anonymous):

prove these both to be right (sinx)/(1-cosx)+ (sinx)/(1+cosx) = 2 csc x - tan2x + sec2x = 1

OpenStudy (anonymous):

i actually have to make the right side match the left side and its tan^2x

OpenStudy (anonymous):

These can be "trivially" true, like "x = x" or usefully true, such as the Pythagorean Theorem's "a2 + b2 = c2" for right triangles. ... sec(x) = 1/cos(x), csc(x) ... sin2(t) + cos2(t) = 1 tan2(t) + 1 = sec2(t) 1 + cot2(t) = csc2(t) ...

OpenStudy (anonymous):

rify each trigonometric equation by substituting identities to match the right hand side ... cot x + 2 tan x + tan3x (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos2x 1 + sec2x sin2x = sec2x [(sin(x))/(1-cos(x))]+[(sin(x))/(1+cos(x))]=2csc(x) - tan2x + sec2x = 1 ...

OpenStudy (anonymous):

yeah i have done a few of these so far but i have gotten stuck. can you help?

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. These are the directions on the sheet

OpenStudy (isaiah.feynman):

So first or second one?

OpenStudy (anonymous):

both if you can

OpenStudy (anonymous):

please

OpenStudy (isaiah.feynman):

You've got to contribute as well.

OpenStudy (anonymous):

I will I just need guidance

OpenStudy (isaiah.feynman):